OpenStudy (anonymous):

can anyone show me the step by step how to solve this ... \[\int\limits_{}^{} \sinh ^{-1}x dx\]

OpenStudy (amistre64):

does inverse of sinh have a nice e format?

OpenStudy (anonymous):

\[\sinh ^{-1}x =\ln (x+\sqrt{x^{2}+1})\]

OpenStudy (amistre64):

hmmm, yeah isnt there some rule similar to sin and cos ??

OpenStudy (amistre64):

i just havent played with the hypers enough to recall it that well; but this might be useful http://www.wolframalpha.com/input/?i=integrate+sinh%5E-1+x

OpenStudy (anonymous):

tq...I think i have to use integration by part... ;)

OpenStudy (amistre64):

integration by parts eh, i was considering something like this \[y =\ln (x+\sqrt{x^{2}+1})\] \[\large e^y =e^{ln (x+\sqrt{x^{2}+1})}\] \[\large e^y =x+\sqrt{x^{2}+1}\] \[\large\int( e^y =x+\sqrt{x^{2}+1})\] \[\large e^y =\int x~dx+\int~\sqrt{x^{2}+1}~dx\] \[\large e^y =\frac12x^2+c+\int~(x^{2}+1)^{1/2}~dx\] u = x^2+1; du = 2x dx \[\large e^y =\frac12x^2+c+\int~\frac22(u)^{1/2}~du\] \[\large e^y =\frac12x^2+c+\frac12\int~2(u)^{1/2}~du\] \[\large e^y =\frac12x^2+c+\frac12\frac23~2(u)^{3/2}+c\] \[\large e^y =\frac12x^2+\frac23(u)^{3/2}+C\] \[\large e^y =\frac12x^2+\frac23(x^2+1)^{3/2}+C\]

OpenStudy (amistre64):

i believe i mismathed it at my attempt to usub

OpenStudy (amistre64):

oh well :) good luck with it tho, ive got to get ;)

OpenStudy (anonymous):

why dont we use...\[u = \ln (x+\sqrt{x^{2}+1})\] \[dv = dx\]